[1 1 0] [1 1 0] [1 \overline{1} 0] [1 \overline{1} 0] (1 \overline{1} 0) ⟨1 2 3 ⟩ ⟨1 2 \overline{3} ⟩ \{\overline{12} \overline{1} 0\} \{\overline{a} \overline{b} c\} \{\overline{1} \overline{1} 21 0\} \{\overline{1} \overline{1} 21 \overline{0.5}\}